Most antenna consisting of a 50 meter long vertical conductor operates over a perfectly conducting ground plane. It is base fed at a frequency of 600 kHz. The radiation resistance of the antenna in ohms is


1
Expert's answer
2021-01-11T11:59:45-0500

Answer

Radiation resistance is given by

"R_{rad}=40 \pi^2 ( \frac{dl}{ \lambda})\\and\\ \\\lambda=\frac{c}{f}\\R_{rad}=40 \pi^2 ( \frac{(dl)(f)}{ c})\\
R_{rad}=40 \pi^2 ( \frac{(50)(600\times10^3)}{ 3\times10^8}) =\frac{2\pi^2}{5} \Omega"